OpenStudy (jiteshmeghwal9):

How many integer solution exist for \(x\) in \(\dfrac{2x^2+2x-30}{x^2+x-12}>3\) :-

OpenStudy (jiteshmeghwal9):

@AravindG @satellite73 @amistre64 plz help :)

OpenStudy (aravindg):

first simplify the inequality

OpenStudy (jiteshmeghwal9):

\[{2(x^2+x-15) \over x^2+x-12}>3\]

OpenStudy (jiteshmeghwal9):

@AravindG

OpenStudy (sirm3d):

\[\frac{2x^2+2x-30}{x^2+x-12}-3>0\\\frac{(2x^2+2x-30)-(3x^2+3x-36)}{x^2+x-12}>0\\\frac{-x^2-x+6}{x^2+x-12}>0\\\frac{(3+x)(2-x)}{(x-3)(x+4)}>0\]  looks like there's no integer solution